Calculate the force of gravity of the husband and wife, the husband has a mass of 85 kg and the wife has a mass of 65 kg. They are 0.75 m apart


1
Expert's answer
2021-01-24T14:23:08-0500

From the Newton's Law of Universal Gravitation, we get:


"F=\dfrac{Gm_1m_2}{r^2},""F=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot85\ kg\cdot65\ kg}{(0.75\ m)^2}=6.5\cdot10^{-7}\ N."